OpenStudy (anonymous):

Sketch the graph of an even function f and state how f(x) and f(-x) are related?

OpenStudy (anonymous):

if \(f\) is even then \(f(x)=f(-x)\) and the graph is symmetric with respect to the \(y\) axis if you want a graph try \(y=|x|\) or \(y=x^2\) as good examples

OpenStudy (puzzler7):

@satellite73 how did you get a score of 100?!?!?!?

OpenStudy (anonymous):

i stole it from a dead mathematician

OpenStudy (anonymous):

oh ok and what if its odd

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

if it is "odd" then \(f(-x)=-f(x)\) like for example \(f(x)=x^3\) or \[f(x)=\frac{x}{x^2-1}\] or even just \[f(x)=x\]

OpenStudy (anonymous):

it also says state how f(x) and f(-x) are related . so for the absolute value graph wut would i say

OpenStudy (anonymous):

@satellite73